# Improving Vision-Language-Action Models via Chain-of-Affordance

### v1

@@ 5.1 Evaluation on Real Robot
Table 1: Experimental Results for Multi-Task Learning. Our method achieved the best performance in both the in-distribution test setup and under visual changes..
@@
DiffusionVLA [43] 2/9 2/9 5/9 5/9 5/9 3/9 6/9 27/63 (42.85%)
CoA 4/9 4/9 6/9 6/9 5/9 4/9 7/9 36/63 (57.14%)
@@
Diffusion Policy [9] $78.3\pm 1.1\%$ $92.5\pm 0.7\%$ $68.3\pm 1.2\%$ $50.5\pm 1.3\%$ $72.4\pm 0.7\%$
ScaleDP [51] $79.1\pm 0.7\%$ $90.4\pm 0.9\%$ $73.6\pm 0.8\%$ $48.4\pm 1.2\%$ $72.9\pm 0.5\%$
Octo [33] $78.9\pm 1.0\%$ $85.7\pm 0.9\%$ $84.6\pm 0.9\%$ $51.1\pm 1.3\%$ $75.1\pm 0.6\%$
OpenVLA [21] $84.7\pm 0.9\%$ $88.4\pm 0.8\%$ $79.2\pm 1.0\%$ $53.7\pm 1.3\%$ $76.5\pm 0.6\%$
CoA $85.3\pm 0.9\%$ $93.1\pm 1.0\%$ $85.8\pm 0.9\%$ $55.0\pm 1.2\%$ $79.8\pm 0.5\%$
Baseline. We compare CoA with Diffusion Policy [9], Octo [33], OpenVLA [21], TinyVLA [44], DiffusionVLA [43]. The latter three methods are all vision-language-action models that achieve state-of-the-art performance in the real world. The DiffusionVLA is the same model our approach is built upon, but trained with vanilla reasoning instead of our proposed chain-of-affordance. To ensure a fair comparison, all models are fine-tuned on the same dataset we use for training our approach. All models are trained with the same number of iterations, and the last checkpoint is used for evaluation. It ensures that no model is cherry-picked for comparison.
Real robot experimental results. The experimental results are presented in Table 1. We assess the performance using two different settings: the in-distribution setup and visual generalization testing. In the in-distribution setup, our method surpasses all SOTA robot foundation models in terms of average success rate. Notably, our method exceeds the performance of OpenVLA by 30.65%, despite having a significantly smaller model size and less pre-training data. Compared to our baseline model, which employs vanilla reasoning, our method achieves a 14.29% increase in accuracy. To assess visual generalization, we further evaluate these advanced models under varying visual conditions. As observations become more complex—such as through the addition of distractors or vibrant backgrounds—the performance gap between CoA and OpenVLA/DiffusionVLA widens. These findings highlight the critical role of embodied, task-specific chain-of-affordance reasoning in enhancing the performance of VLA models.
@@ 5.2 Evaluation on Simulation
In this section, we examine the performance of CoA on LIBERO [27]. LIBERO is a robot learning benchmark comprising over 130 language-conditioned manipulation tasks. We follow the setting as in OpenVLA [21] open-sourced code and test on four task suites: LIBERO-Spatial, LIBERO-Goal, LIBERO-Object, and LIBERO-Long. The detailed experimental setup are in the Appendix.
Simulation experimental results. We compare with the Diffusion Policy [9], ScaleDP [51], Octo [33], and OpenVLA [21]. The experimental results of LIBERO are presented in Table 2. Our findings indicate that CoA consistently achieves superior performance across all evaluated settings, securing the highest success rate among the methods tested. Specifically, CoA achieves an overall success rate of 79.8%, outperforming OpenVLA, the previous best-performing method, by a margin of 3.3%. This improvement demonstrates the effectiveness of our approach in simulation. Compared to pre-trained robotic models like Octo and OpenVLA, CoA demonstrates non-trivial gains, showing its effectiveness in leveraging prior knowledge while adapting to diverse task requirements. Additionally, when benchmarked against train-from-scratch methods such as Diffusion Policy and ScaleDP, CoA exhibits substantial improvements, underscoring the advantages of our approach.
Figure 3: Spatial Affordance for CoA. CoA can identify free space for object placement..
Figure 4: Movement Generalization for CoA. CoA can avoid obstacles and operate safely.
@@ 5.3 More Experiments
Spatial affordance enables safe placement. We evaluate the effectiveness of our spatial affordance approach in the PlaceBread task, as illustrated in Figure 3. In this task, the robot is presented with a plate on which three distinct objects are already placed, and it is instructed to add a piece of bread onto the plate. Our method successfully identifies open areas on the plate, allowing it to accurately position the bread without interference, thereby enabling CoA to complete all three task scenarios with precision. In contrast, alternative methods, such as OpenVLA and DiffusionVLA, succeed in only one scenario each, failing to generalize across all spatial configurations. This experiment underscores the crucial role of spatial affordance in enhancing the model’s ability to recognize and utilize available space, ultimately improving task completion accuracy and reliability across varied setups.
Obstacle avoidance. Collision avoidance is essential for safe and effective physical interactions, as improper maneuvers can lead to significant damage or even catastrophic outcomes. We evaluated CoA’s obstacle avoidance capabilities in two specific tasks, illustrated in Figure 4. In the first task, we positioned a vase near the center of the task area to observe whether CoA could successfully navigate around the vase to retrieve a piece of paper trash. In the second task, we introduced a series of obstacles on a table, rearranging them in different configurations to assess the robot’s adaptability in maneuvering through varied layouts to complete the task. Our approach successfully completed all three scenarios, demonstrating robust collision avoidance and spatial adaptability. In contrast, OpenVLA failed to complete any of the tasks, while DiffusionVLA succeeded in only one scenario. These results highlight the critical role of integrating movement affordance into the model’s reasoning process, enhancing its ability to navigate complex environments and complete tasks with precision and safety.
Generalization to unseen object pose. We assessed CoA’s ability to generalize to previously unseen object orientations, as illustrated in Figure 5. Our evaluation focused on two objects: a hammer and a teapot. In the training phase, both objects were consistently presented with their handles oriented vertically relative to the robot. To test the model’s generalization capabilities, we introduced novel poses that were absent from the training data, challenging CoA to grasp these objects in unfamiliar orientations. We observed that CoA successfully managed most scenarios, demonstrating a remarkable ability to adapt to new object poses even without explicit training on these orientations. In contrast, OpenVLA succeeded only in the simplest cases, struggling with more complex orientations. However, when the objects were positioned horizontally relative to the robot, all models, including CoA, were unsuccessful in achieving a stable grasp. Despite this limitation, our grasp affordance approach shows promising results, enabling CoA to handle a wide range of novel object poses.
@@ 6 Conclusion
Explicit reasoning is essential for language models to handle complex tasks. In this work, we design a reasoning-aware foundation model for robotics, focusing on various affordances: object, grasp, spatial, and movement. These affordances form an interdependent chain: the robot identifies the target object and location, determines how to grasp it, decides where to place it, and navigates accordingly. By structuring this chain of affordances as intermediate language and image outputs and feeding this reasoning into the policy model, our Chain-of-Affordances (CoA) model outperforms baselines on real-world robotic tasks. CoA models also generalize well to complex environments, tackling challenges like grasping in unfamiliar orientations, avoiding obstacles, and spatial generalization. Our approach provides a novel perspective on designing reasoning chains to enhance embodied control in real-world settings.
@@ 8 Evaluation Tasks
9 Details for Real Robot Experiments
@@
10 Details for LIBERO Simulation
@@
Each suite includes 10 distinct tasks with 50 demonstrations per task. Each task suite emphasizes unique challenges in imitation learning: LIBERO-Goal features tasks with similar object categories but different goals. LIBERO-Spatial requires policies to adapt to varying spatial arrangements of the same objects. LIBERO-Object keeps the layout consistent while changing the objects. During experimentation, our method uses a static camera, and a wrist-mounted camera all methods are evaluated across 1500 trials in total. We filter out the failure data and increase the image resolution to $224\times 224$. The affordance data is generated using our proposed pipeline for data in LIBERO. In Table 2, we directly cite the results of Diffusion Policy, Octo, and OpenVLA from OpenVLA’s paper. Therefore, to ensure all methods are evaluated fairly, we evaluated our methods across 500 trials for each task suite, and the reported performance is the average success rate over three random seeds. We use the same test data as in OpenVLA. For the baseline ScaleDP, except for using all camera views, all other implementations kept the same.

### v2

@@ 4.2 Formatting of Visual-Textual Chain-of-Affordance
Table 1: Experimental results for multi-task learning. Our method achieved the best performance in both the in-distribution test setup and under visual changes.
@@
DiffusionVLA [46] 2/9 2/9 5/9 5/9 5/9 3/9 6/9 28/63 (44.44%)
CoA-VLA 4/9 4/9 6/9 6/9 5/9 4/9 7/9 36/63 (57.14%)
4.3 Generating Chain-of-Affordance Data
Our pipeline begins with GPT-4o [38], which generates a detailed description of the scene and identifies relevant entities from the language instructions. This allows us to create contextually rich, entity-specific affordances tailored to the task at hand. Using these entities, we leverage Grounding DINOv2 [32] and SAM [23, 40] to produce bounding boxes around each identified object within the scene. SAM initially provides masks, which we convert into bounding boxes, and we finalize these by taking out the intersection over union (IoU) between the outputs of Grounding DINOv2 and SAM. This IoU-based refinement ensures that bounding boxes are accurately aligned with object contours. At this stage, we also capture the gripper’s position for subsequent affordance calculations.To represent spatial affordance, we integrate RoboPoint [51], a state-of-the-art model that predicts spatial affordances directly within the image. Additionally, we prompt GPT-4o to annotate spatial points based on the scene context. The spatial predictions from RoboPoint and GPT-4o are then combined, after which we cluster these points to form a coherent representation, eliminating any outliers to maintain accuracy. This process captures spatial affordances that are not only visually aligned but contextually relevant.For capturing movement trajectories, we employ CoTracker [19, 20], an advanced transformer-based tracking model. CoTracker enables us to follow the robot gripper’s path, recording its movement through the scene and gathering essential trajectory data. This movement data provides insight into how the robot interacts dynamically with the environment, adding temporal dimensions to our affordance representation. The combination of spatial affordances, grasping points, and movement trajectories enables us to model a rich, multi-faceted affordance landscape for each scenario. This automated, tool-assisted pipeline produces a detailed and diverse set of affordance annotations, significantly reducing the need for manual labeling.
@@ 4.3 Generating Chain-of-Affordance Data
[…]
5 Experiments
The objective of our experimental evaluations is to assess the efficacy of CoA-VLA as a reasoning framework that augments the policy-learning capabilities of a baseline robot foundation model. Specifically, we seek to address the following questions: 1) Does the VLA model demonstrate improved performance with the integration of CoA in real-world experiments? 2) What is the performance of CoA-VLA in simulation benchmarks? 3) How effectively does CoA-VLA generalize to challenging scenarios? 4) How important is our proposed visual-textual affordance approach, compared to the vanilla VLA? We will give a detailed analysis in this section. Note that some experiments are presented in the Appendix due to limited space.
5.1 Evaluation on Real Robot
Experimental setup. We conduct evaluations using the 6-DoF Franka robot arm, a widely adopted setup for assessing generalizable robot policies. Our setup includes two third-person cameras (ZED cameras) positioned on either side of the robot arm, as well as an egocentric camera (Realsense D435i) mounted on the wrist. We design seven challenging tasks, such as pouring tea and cleaning the table, incorporating both short-horizon and long-horizon tasks to test the adaptability of our approach. We illustrate our robot setup and tasks with examples in Figure 3. Detailed descriptions of each task and the experimental setup, and our ablation experiments are provided in the Appendix.
Train data. We use the Droid dataset [21] as an external data source, filtering out samples without language annotations, leaving 39K trajectories. Our pipeline generates synthetic chain-of-affordance data for pre-training. The model is then post-trained on 692 trajectories across seven tasks.
Baseline. We compare CoA-VLA with Diffusion Policy [10], Octo [36], OpenVLA [22], TinyVLA [47], DiffusionVLA [46]. The latter three methods are all vision-language-action models that achieve state-of-the-art performance in the real world. Notably, the DiffusionVLA is the same model our approach is built upon, but trained with vanilla reasoning instead of our proposed chain-of-affordance. To ensure a fair comparison, all models are fine-tuned on the same dataset we use for training our approach. All models are trained with the same number of iterations, and the last checkpoint is used for evaluation. It ensures that no model is cherry-picked for comparison.
Real robot experimental results. The experimental results are presented in Table 1. We assess the performance using two different settings: the in-distribution setup and visual generalization testing. In the in-distribution setup, our method surpasses all SOTA robot foundation models in terms of average success rate. Notably, our method exceeds the performance of OpenVLA by 30.65%, despite having a significantly smaller model size and less pre-training data. Compared to our baseline model, which employs vanilla reasoning, our method achieves a 14.29% increase in accuracy. To assess visual generalization, we further evaluate these advanced models under varying visual conditions. As observations become more complex—such as through the addition of distractors or vibrant backgrounds—the performance gap between CoA-VLA and OpenVLA/DiffusionVLA widens. These findings highlight the critical role of embodied, task-specific chain-of-affordance reasoning in enhancing the performance of VLA models.
@@ 5.2 Evaluation on Simulation
In this section, we examine the performance of CoA-VLA on LIBERO [30]. LIBERO is a robot learning benchmark comprising over 130 language-conditioned manipulation tasks. We follow the setting as in OpenVLA [22] open-sourced code and test on four task suites: LIBERO-Spatial, LIBERO-Goal, LIBERO-Object, and LIBERO-Long. The detailed experimental setup is in the Appendix.
Simulation experimental results. We compare with the Diffusion Policy [10], ScaleDP [56], Octo [36], and OpenVLA [22]. The experimental results of LIBERO are presented in Table 2. Our findings indicate that CoA-VLA consistently achieves superior performance across all evaluated settings, securing the highest success rate among the methods tested. Specifically, CoA-VLA achieves an overall success rate of 79.8%, outperforming OpenVLA, the previous best-performing method, by a margin of 3.3%. This improvement demonstrates the effectiveness of our approach in simulation. Compared to pre-trained robotic models like Octo and OpenVLA, CoA-VLA demonstrates non-trivial gains, showing its effectiveness in leveraging prior knowledge while adapting to diverse task requirements. Additionally, when benchmarked against train-from-scratch methods such as Diffusion Policy and ScaleDP, CoA-VLA exhibits substantial improvements, underscoring the advantages of our approach.
Figure 4: Spatial affordance for CoA-VLA. CoA-VLA  can identify free space for object placement..
Figure 5: Movement generalization for CoA-VLA. CoA-VLA can avoid obstacles and operate safely.
@@ 5.3 More Experiments
Spatial affordance enables safe placement. We evaluate the effectiveness of our spatial affordance approach in the PlaceBread task, as illustrated in Figure 4. In this task, the robot is presented with a plate on which three distinct objects are already placed, and it is instructed to add a piece of bread onto the plate. Our method successfully identifies open areas on the plate, allowing it to accurately position the bread without interference, thereby enabling CoA-VLA to complete all three task scenarios with precision. In contrast, alternative methods, such as OpenVLA and DiffusionVLA, succeed in only one scenario each, failing to generalize across all spatial configurations. This experiment underscores the crucial role of spatial affordance in enhancing the model’s ability to recognize and utilize available space, ultimately improving task completion accuracy and reliability across varied setups.
Obstacle avoidance. Collision avoidance is essential for safe and effective physical interactions, as improper maneuvers can lead to significant damage or even catastrophic outcomes. We evaluated CoA-VLA’s obstacle avoidance capabilities in two specific tasks, illustrated in Figure 5. In the first task, we positioned a vase near the center of the task area to observe whether CoA-VLA could successfully navigate around the vase to retrieve a piece of paper trash. In the second task, we introduced a series of obstacles on a table, rearranging them in different configurations to assess the robot’s adaptability in maneuvering through varied layouts to complete the task. Our approach successfully completed all three scenarios, demonstrating robust collision avoidance and spatial adaptability. In contrast, OpenVLA failed to complete any of the tasks, while DiVLA succeeded in only one scenario. These results highlight the critical role of integrating movement affordance into the model’s reasoning process, enhancing its ability to navigate complex environments and complete tasks with precision and safety.
@@ 6 Conclusion
Explicit reasoning is essential for language models to handle complex tasks. In this work, we design a reasoning-aware foundation model for robotics, focusing on various affordances: object, grasp, spatial, and movement. These affordances form an interdependent chain: the robot identifies the target object and location, determines how to grasp it, decides where to place it, and navigates accordingly. By structuring this chain of affordances as intermediate language and image outputs and feeding this reasoning into the policy model, our Chain-of-Affordances (CoA-VLA) model outperforms baselines on real-world robotic tasks.CoA-VLA models also generalize well to complex environments, tackling challenges like grasping in unfamiliar orientations, avoiding obstacles, and spatial generalization. Our approach provides a novel perspective on designing reasoning chains to enhance embodied control.
Acknowledegments
This work is supported by the National Science Foundation of China (12471501), and the Sci-Tech Innovation Initiative by the Science and Technology Commission of Shanghai Municipality (24ZR1419000).
@@ 6.2 Evaluation Tasks
6.3 Details for Real Robot Experiments
@@
Table 4: Ablation study on visual affordance and textual affordance. Our experiments demonstrate that both affordance are important for VLA.
LIBERO-Spatial LIBERO-Object LIBERO-Goal LIBERO-Long Average
[…]
Table 5: Ablation study on dynamic affordance selection. Removing dynamic affordance selection causes introduction of redundant affordance into the learning process, which cause the model to perform even worse than the baseline without it.
LIBERO-Spatial LIBERO-Object LIBERO-Goal LIBERO-Long Average Inference
Method / Task Success Rate ($\uparrow$) Success Rate ($\uparrow$) Success Rate ($\uparrow$) Success Rate ($\uparrow$) Success Rate ($\uparrow$) Speed
[…]
6.4 Details for LIBERO Simulation
@@
Each suite includes 10 distinct tasks with 50 demonstrations per task. Each task suite emphasizes unique challenges in imitation learning: LIBERO-Goal features tasks with similar object categories but different goals. LIBERO-Spatial requires policies to adapt to varying spatial arrangements of the same objects. LIBERO-Object keeps the layout consistent while changing the objects. During experimentation, our method uses a static camera, and a wrist-mounted camera all methods are evaluated across 1500 trials in total. We filter out the failure data and increase the image resolution to $\times 224$ 224. The affordance data is generated using our proposed pipeline for data in LIBERO. In Table 2, we directly cite the results of Diffusion Policy, Octo, and OpenVLA from OpenVLA’s paper. Therefore, to ensure all methods are evaluated fairly, we evaluated our methods across 500 trials for each task suite, and the reported performance is the average success rate over three random seeds. We use the same test data as in OpenVLA. For the baseline ScaleDP, except for using all camera views, all other implementations kept the same.
7 More Experiments
7.1 Ablation Study on Visual-Textual Affordance
Our primary contribution lies in the introduction of textual affordances and visual affordances, paired with a novel visual-textual co-injection module designed to synergistically integrate these modalities into policy learning. To validate their individual and combined efficacy, we conduct a systematic ablation study (Table 4) on the LIBERO robotic task benchmark. Our key finding is that both textual and visual affordances are critical to model performance. Removing either modality leads to significant degradation in task success rates. While both modalities contribute uniquely, textual affordances exhibit stronger influence on policy optimization. We hypothesize that this stems from language’s inherent capacity to encode task-specific semantics (e.g., ”pour-able” or ”graspable”), which provides clearer optimization signals compared to visual features that require implicit spatial grounding. These results underscore the importance of our co-injection module, which dynamically balances and fuses multimodal affordances to maximize policy robustness in diverse environments.
7.2 Ablation Study on Dynamic Affordance Selection
Utilizing all affordances can be computationally expensive and time-consuming. Therefore, we introduce a dynamic affordance selection mechanism. This approach focuses on selectively utilizing only the most relevant affordances at each time step. As demonstrated in Table 5, our method outperforms a baseline model that employs all affordances indiscriminately. Surprisingly, using all affordances results in a lower average success rate compared to our dynamic selection approach. We hypothesize that the irrelevant affordances introduce noise during the optimization process, hindering the model’s learning ability. To further analyze the impact of dynamic selection, we measured inference speed on an Nvidia 3090 GPU. We averaged the running time over all tasks, with each task measured across 5 trials. Our results show that utilizing all affordances significantly impacts inference speed, causing the model to run 6 times slower than our proposed method. This highlights the substantial efficiency gains achieved through dynamic affordance selection
7.3 Generalization to Unseen Object Pose
We assessed CoA-VLA’s ability to generalize to previously unseen object orientations, as illustrated in Figure 8. Our evaluation focused on two objects: a hammer and a teapot. In the training phase, both objects were consistently presented with their handles oriented vertically relative to the robot. To test the model’s generalization capabilities, we introduced novel poses that were absent from the training data, challenging CoA-VLA to grasp these objects in unfamiliar orientations. We observed that CoA-VLA successfully managed most scenarios, demonstrating a remarkable ability to adapt to new object poses even without explicit training on these orientations. In contrast, OpenVLA succeeded only in the simplest cases, struggling with more complex orientations. However, when the objects were positioned horizontally relative to the robot, all models, including CoA-VLA, were unsuccessful in achieving a stable grasp. Despite this limitation, our grasp affordance approach shows promising results, enabling CoA-VLA to handle a wide range of novel object poses.